OpenStudy (anonymous):

A buys a camera for $1500 and sells it to B at a profit of 20%.B sells it to C at a loss of 20%.What does C pay for the camera?

OpenStudy (anonymous):

First, look at what B paid for the camera. If A made a 20% profit, he sold the camera for 1500*1.2=1800. Now since B took a 20% loss, C bought the camera for 1800*.8=?

OpenStudy (asadkarim7):

$ 1500 x 20/ 100 =$300 A 's Selling price 1500 +300 = 1800 1800 x 20/100 = $360 B's selling price = 1800 -360 = 1440 C paid $1440 for the camera